Given that P(A) = 0.5, P(B) = 0.6, and P(A and B) = 0.10, determine P(A|B)
Choices are 0.30, 0.17, 0.11, 0.20

Answer by jim_thompson5910(35256)

P(A|B) = P(A intersect B)/P(B)

P(A|B) = P(A and B)/P(B)

P(A|B) = 0.1/0.6

P(A|B) = 0.1667

which rounds to P(A|B) = 0.17

So the answer is choice B
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------